OpenStudy (anonymous):

A potato is taken from the oven, its temperature having reached 350 degrees. After sitting on a plate in a 70 degree room for 12 minutes, its temperature has dropped to 250 degrees. If Ted will only eat his potato if its between 100 and 150 degrees, will he eat the potato if he arrives home for dinner 20 minutes after the potato came out of the oven? Show all your work and EXPLAIN YOUR REASONING. RATIONALIZE YOUR ANSWER! Assume Newton's Law Of Cooling, which says that the difference between an object's temperature and the ambient temperature is an exponential function of time.

OpenStudy (anonymous):

rawr.... please help me, i need to finish this TONIGHT

OpenStudy (anonymous):

only information i have for the equation \[y=ab ^{x}\]

OpenStudy (agreene):

does it define a or b ?

OpenStudy (anonymous):

im not sure

OpenStudy (anonymous):

dumb teacher gave this to us without explaining, its already a late assignment and its graded

OpenStudy (agreene):

Is this a differential equations class? Because if I'm remembering correctly--Newtons Law of Cooling is a DiffEQ

OpenStudy (anonymous):

-_- im in Pre-calculus

OpenStudy (anonymous):

ive never even heard of differential equations before

OpenStudy (anonymous):

our teacher is nuts

OpenStudy (agreene):

This is a really annoying question... lol

OpenStudy (anonymous):

you see my problem? and ive never seen anything like this before

OpenStudy (agreene):

Ive seen lots of things similar to this--but usually they give more information as to wtf they are talking about, lol. And I know the real Newton's Law of Cooling--and it isn't useful at all for this

OpenStudy (anonymous):

we are only barely starting to learn logarithms

OpenStudy (anonymous):

ergh only thing i could figure out, is that b=280

OpenStudy (anonymous):

my bad, a=280 not b

OpenStudy (agreene):

if you make that the case you end up with a really messy eqn... y=a*b^x y=280*b^x up in the question it says its an exponential function of time... meaning 250=280*b^12 you are gonna get something really disgusting for b.

OpenStudy (anonymous):

i think its something like .96

OpenStudy (agreene):

of course, if you make it b... it isn't quite as bad 250=a*280^12 a is just gonna be a really small number

OpenStudy (agreene):

250=280*b^12 b=-0.9906003 or b=0.9906003 but showing your work is gonna be lame.

OpenStudy (anonymous):

rawr, i found that a was a=280 like this difference in temperature from intial intial temp=350 time=0 average temp=70 diiference in temp =280 x would be time and y would be the diference in temp plugin ---> 280=ab^0 ---> (becomes) 280=a cuz anynumber with an exponent of 0 is = to 1

OpenStudy (anonymous):

whatever, i give up, my whole thing is messed up ill just get an F

OpenStudy (anonymous):

too much of a pain